Session: Sep 2000 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Tue Sep 19 10:33:51 METDST 2000) Success
No known problems.

Onsala (A. Polatidis at Tue Sep 19 11:32:43 METDST 2000) Success
We are observing LINEAR polarization fed to both IF1 and IF2

Noto (C.Trigilio at Wed Sep 20 14:13:32 METDST 2000) Success

Effelsberg (A. Kraus at Thu Sep 21 14:48:04 METDST 2000) Minor failures
LCP and RCP are exchanged! The receivers RCP-channel has much higher system temperature than LCP. More details on the 5cm-session page.

Jodrell_Bank (Alastair Gunn at Tue Sep 26 06:28:58 METDST 2000) Minor failures
Severe problems. FS underwent frequent crashes, 12 in all. Also, faulty RF component caused comb structure in band. This was corrected by lowering LO from correct value (295 MHz) to incorrect value (275 MHz) and back again. Data should be fine during periods of 295 MHz LO. I estimate that at least half of data should be good, it just needs to be found by the correlator...